# REDS: Resource-Efficient Deep Subnetworks for Dynamic Resource Constraints

### v2

@@ 3.3 Iterative knapsack problem
This section describes the novel method to design REDS subnetwork structure by formulating and solving a variant of an iterative knapsack problem. The discussion in the main paper is limited to the simplest case, while the generalization to more complex model architectures, handling dependencies between layers and the theoretical analysis of the problem are moved to Appendix A.1 and Appendix A.2 for presentation clarity.
@@ 3.3.1 Problem formulation
Given a pre-trained model, REDS identifies how to best slice the weight tensors by formulating the problem as an iterative knapsack problem with $k$ stages: the items included in a knapsack with capacity $c$ have to be included in all later stages, i.e., knapsacks with larger capacities (Della Croce et al., 2019; Faenza et al., 2023). The items correspond to all the computational units that compose the model encoder architecture. Given a list of MACs and peak memory usage values REDS solves an iterative knapsack problem with as many stages as the predefined number of subnetworks. We give two heuristic algorithms that solve the corresponding iterative knapsack problem and hence find subnetwork architectures, i.e., slicing points for each layer and each subnetwork, that satisfy these capacity constraints while maximizing $I$ to keep the most important units in each subnetwork. Both heuristics are based on solving several knapsack problems formulated as integer programs (Perron and Furnon, ) which we then solve using a mixed integer programming solver (Bixby, 2007). Due to permutation invariance, the items are stored and passed to the solver in descending order of their importance scores. This property makes the solver select items in descending order, thereby creating contiguous active and inactive units for the weight tensors in each subnetwork. Given $C_{MACs}$ as the maximum number of MACs in a subnetwork and $C_{PeakMem}$ as the maximum peak memory usage for the model, i.e., the maximum size of each layer activation maps (Liberis and Lane, 2023), a single stage knapsack problem is formulated as follows:
@@
$\sum_{l=1}^{L}\left(\sum_{i=1}^{u_{l}}x_{il}\cdot PeakMem_{il}\right)\leq C_{PeakMem}\,,$
$x_{il}\in\{0,1\},\quad\forall\,l\in\{1,\dots,L\},\;\;i\in\{1,\dots,u_{l}\},$
$I_{i1}\geq I_{i2}\geq\dots I_{il},$
where $x_{il}$ is a binary decision variable taking value 1 if an item $i$ from layer $l$ is selected and 0 otherwise; $I_{il}$ is the importance score of item $i$ from layer $l$; $MACs_{il}$ is the number of MACs of item $i$ from layer $l$; $PeakMem_{il}$ is the number of byte of the activation map of item $i$ from layer $l$; $L$ is the number of layers in the model and $u_{l}$ is the number of items in layer $l$. The above knapsack problem formulation does not account for dependencies between layers. Its generalization and adaptation for depthwise separable convolutions (Howard et al., 2017), for simplicity considering only $MACs$ as a constraint, is moved to Appendix A.1 for clarity of presentation.
3.3.2 Bottom-up (BU) and top-down (TD) heuristics
REDS examines two heuristics for solving the iterative knapsack problem named bottom-up (BU) and top-down (TD). The former iteratively calculates the subnetwork architectures by considering the tightest constraints for the smallest subnetwork first. Once a solution is found by the knapsack solver, these units are frozen, i.e., they are now part of all nested subnetworks, and the second smallest subnetwork is being computed by the solver. The latter top-down method determines solutions by considering the weakest constraints first and then iteratively searching the architectures for increasingly smaller subnetworks. Related works, including DRESS and NestDNN, use a variant of the top-down approach. Our theoretical analysis of the iterative knapsack problem (based on the classical 0-1 knapsack problem) in Appendix A.2 shows that the bottom-up approach promises a better worst-case performance. In particular, we prove that the solution found by the two-stage bottom-up iterative knapsack heuristic is not worse than $\frac{2}{3}\cdot Opt$, where $Opt$ is the optimal solution of the knapsack with larger capacity, and that this bound is tight. We then show that the tight bound for the top-down iterative knapsack is $\frac{1}{2}\cdot Opt$, where $Opt$ in this case is the optimal solution of the knapsack with smaller capacity. Since our generalized problem suited for depthwise separable convolutional architectures has the classical 0-1 knapsack as its core problem, we believe that a similar result is valid for this case as well.
@@ 3.4 REDS fine-tuning
Figure 3: REDS fine-tuning. The parameters ${\pi_{i}}_{i=1}^{N}$ ensure the contribution of individual models to the loss aligns with the fraction of the shared weights (Qu et al., 2022).
@@
Fine-tuning models after pruning is a common practice to recover model accuracy. The REDS subnetworks found by the heuristics algorithm are fine-tuned simultaneously to recover their full accuracy. We first fine-tune the encoder’s weights followed by the batch normalization layers (Qu et al., 2022). For each subnetwork we slice each layer to construct a subnetwork architecture by storing only the slicing point corresponding to the number of active computational units (see Fig. 3). During fine-tuning, the weights of each subnetwork $i$ are reused by all lower-capacity subnetworks $\{i+k\}_{k=1}^{N}$ by dynamically creating a tensor slice for each layer, i.e., a tensor view object that points to the original weight tensor. We balance the contribution of the loss of each individual model with parameters $\{\pi_{i}\}_{i=1}^{N}$ equal to the percentage of weights used in the encoder part of the model by the subnetwork $i$ (Qu et al., 2022).

### v3

@@ 3.4 Iterative knapsack problem
This section describes the novel method to design REDS subnetwork structure by formulating and solving a variant of an iterative knapsack problem. The theoretical analysis is moved to the Appendix A for presentation clarity.
@@ 3.4.1 Problem formulation
Given a pre-trained model, REDS identifies how to best slice the weight tensors by formulating the problem as an iterative knapsack problem with $k$ stages: the items included in a knapsack with capacity $c$ have to be included in all later stages, i.e., knapsacks with larger capacities (Della Croce et al., 2019; Faenza et al., 2023). The items correspond to all the computational units that compose the model encoder architecture, i.e., filters and kernels. Given $C_{MACs}$ as the maximum number of MACs, for a subnetwork $s$, a single stage knapsack problem is formulated as follows:
@@
$x_{il}\in\{0,1\},\quad\forall\,l\in\{1,\dots,L\},\;\;i\in\{1,\dots,u_{l}\}.$
where $x_{il}$ is a binary decision variable taking value one if an item $i$ from layer $l$ is selected and zero otherwise; $I_{il}$ is the importance score of item $i$ from layer $l$; $MACs_{il}$ is the number of MACs of item $i$ from layer $l$; $L$ is the number of encoder layers in the model and $u_{l}$ is the number of items in layer $l$. As an extension of this model we added the constraint to each knapsack problem that one unit per layer must be taken. This guarantees that the subnetworks extraction will not degenerate into a trivial solution, i.e., pruning whole layers.
3.4.2 Knapsack for depth-wise convolutions
We define a novel generalized knapsack framework for depth-wise convolutions (Howard et al., 2017). Given a $m\times n$ input matrix $A$, the first layer $L_{0}$ is a standard convolutional layer (cf. Fig. 2) applied to $A$ that consists of $N_{0}$ filters and hence produces $N_{0}$ output channels. It is followed by depth-wise convolutional blocks $B_{1},\ldots,B_{d}$, where each block $i$ is built by a depth-wise layer $B_{i}^{D}$ and a point-wise layer $B_{i}^{P}$ for $i=1,\ldots,d$. A depth-wise layer applies one filter on each of the $N_{i-1}$ input channels. The point-wise layer $B_{i}^{P}$ then applies $N_{i}$ filters, where the kernel number for each filter has to be equal to the number of filters of the previous layer. If we want to reduce the size of the network, we can decide on how many filters we use at $L_{0}$ and at each $B_{i}^{P}$. E.g., if we choose $k$ filters at $L_{0}$, the layer $B_{1}^{D}$ will have $k$ filters and one filter of $B_{1}^{P}$ will have $k$ as kernels number. Integer programming allows structural pruning of a neural network, i.e., choosing an optimal number of filters of $L_{0}$ and of $B_{i}^{P}$ to maximize performance of the network obeying a constraint on the number of MACs.
Integer decision variables $x_{0}$ are used for the number of filters at $L_{0}$, and $x_{i}$ for the number of filters at each $B_{i}^{P}$. Other decision variables control whether a unit is used by a subnetwork or not. For $L_{0}$ we introduce $N_{0}$ binary variables $y_{1},\ldots,y_{N_{0}}$, and for every block $B_{i}$, $i=1,\ldots,d$, binary variables $f_{k}^{i}$ and $g_{kt}^{i}$, $k\in\{1,\ldots,N_{i}\}$ and $t\in\{1,\ldots,N_{i-1}\}$, $f_{k}^{i}$ to indicate whether a filter $k$ is used in $B_{i}^{P}$, and $g_{kt}^{i}$ if filter $k$ with kernel $t$ of $B_{i}^{P}$ is used. For $B_{i}^{D}$ binary variables $d_{t}^{i}$ decide if a depth-wise filter $t\in\{1,\ldots,N_{i-1}\}$ is used. $P^{1}$ is the importance score of a standard convolution filter $i$ in the first layer and $W_{1}$ its number of MACs. For each depth-wise point-wise block $i$, $P^{i}_{t}$ is the importance score for the depth-wise filter $t$, $W_{2}^{i}$ its number of MACs, $P^{i}_{kt}$ is the importance score of the corresponding kernel $k$ of the point-wise filter $t$ in the subsequent layer $N_{i}$ and $W_{3}^{i}$ its number of MACs.
The knapsack formulation for a single stage problem of depth-wise convolutions is described by the following ILP-model, where all variables are binary.
  • (3)
    Maximization of the total importance score of the chosen architecture.
  • (9)
    Satisfaction of constraint $C$ (i.e., MACs).
  • (9)
    The number of filters chosen in the first convolution layer is represented by $x_{0}$, and the number of filters chosen in the depth-wise layer $i$ must match the number of filters picked in the previous layer $i-1$.
  • (9)
    The number of the point-wise filters chosen in layer $i$ is set to $x_{i}$, and the point-wise filters are chosen starting at the first point-wise filter to be contiguous in memory.
  • (9)
    If kernel $t$ of filter $k$ is chosen then the whole filter $k$ is chosen. A point-wise filter $k$ is chosen only if one of its kernels is chosen.
  • (9)
    The number of kernels in the filter $k$ of point-wise layer $i$ must be less or equal to the number of filters taken in the previous depth-wise layer.
  • (9)
    If filter $k$ in layer $i$ is chosen then constraints (9) and (9) ensure that the number of kernels $t$ of filter $k$ in layer $i$ equals the number of point-wise filters in the previous block. If filter $k$ in layer $i$ is not chosen, constraints (9) and (9) imply that all kernels $t$ of filter $k$ at layer $i$ are zero ($N_{i-1}$ is an upper bound on $x_{i-1}$).
$\max$ $\sum_{i=1}^{N_{0}}y_{i}\cdot P^{1}+\sum_{i=1}^{d}\sum_{t=1}^{N_{i-1}}\left(d_{t}^{i}\cdot P^{i}_{t}+\sum_{k=1}^{N_{i}}g_{kt}^{i}\cdot P^{i}_{kt}\right)$ (3)
s.t. $\sum_{i=1}^{N_{0}}y_{i}\!\cdot\!W_{1}+\!\!\sum_{i=1}^{d}\!\sum_{t=1}^{N_{i-1}}\!\!\left(d_{t}^{i}\!\cdot\!W_{2}^{i}+\!\!\sum_{k=1}^{N_{i}}g_{kt}^{i}\!\cdot\!W_{3}^{i}\!\right)\!\leq C$ (9)
$\sum_{i=1}^{N_{0}}y_{i}=x_{0}\;\;\;\;\;\text{and}\;\;\;\;\sum_{t=1}^{N_{i-1}}d_{t}^{i}=x_{i-1}\enspace\forall i$
$\sum_{k=1}^{N_{i}}f_{k}^{i}=x_{i}\enspace\forall i\;\;\;\;\;\text{and}\;\;\;\;f_{k}^{i}\geq f_{k+1}^{i}\enspace\forall i,k$
$g_{kt}^{i}\leq f_{k}^{i}\enspace\forall i,k,t\;\;\;\;\;\text{and}\;\;\;\;f_{k}^{i}\leq\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\enspace\forall i,k$
$\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\leq x_{i-1}\enspace\forall i,k$
$\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\geq x_{i-1}-(1-f_{k}^{i})\cdot N_{i-1}\enspace\forall i,k$
3.4.3 Knapsack for Vision Transformers
We define a generalized knapsack framework for Vision Transformer (Dosovitskiy et al., 2020). Given a $c\times h\times w$ input matrix $A$, the first block $P_{0}$ is a patch embedding block using a convolutional layer to linearly project $A$ into a sequence of patches. The patch embedding convolutional layer consists of convolutional filters whose embedding dimension corresponds to the $N_{0}$ attention heads used in the subsequent ViT layers. The patch embedding block is followed by $B_{1},\ldots,B_{d}$ blocks, where each block $t$ is composed by a Multi-Head Self-Attention (MSA) layer $B_{t}^{MSA}$ and two multi layer perceptron layers (i.e., linear layers) $B_{t}^{\mathrm{mlp1}}$ and $B_{t}^{\mathrm{mlp2}}$ for $t=1,\ldots,d$. Each MSA layer is composed of a $MSA_{qkv}$ linear layer consisting of the concatenation of the query, key and value matrices with their corresponding attention heads. It is followed by a linear projection layer $MSA_{proj}$. Residual connections enforce that the input and output latent patch vectors of each MSA and MLP layer have the same dimension. Structural pruning of a ViT model via integer programming is formulated as choosing the optimal number of attention heads’ convolutional filters $N_{0}$, that reduce each patch dimension of sixty-four for each head’s convolutional filters pruned (Haberer et al., 2024), and the number of neurons of each $B_{i}^{\mathrm{mlp1}}$ to maximize the performance of the network obeying a constraint (i.e., number of MACs).
Integer decision variables $x_{0}$ are used for the number of filters at $P_{0}$ and variables $n_{\mathrm{mlp1_{t}}}$ for the number of neurons at each $B_{t}^{\mathrm{mlp1}}$. For $P_{0}$ we introduce $N_{0}$ binary variables $y_{1},\ldots,y_{N_{0}}$, for every block layer $B_{t}^{\mathrm{mlp1}}$, $t=1,\ldots,d$, binary variables $f_{ik}^{t}$ and for every block layer $B_{t}^{\mathrm{mlp2}}$ binary variables $g_{ki}^{t}$ where $i\in\{1,\ldots,N_{0}\}$ and $k\in\{1,\ldots,N_{i}\}$, $f_{ik}^{t}$ to indicate whether weights $k$ in the neuron $i$ of $B_{t}^{\mathrm{mlp1}}$ are used, and $g_{ki}^{t}$ to indicate whether weights $i$ in the neuron $k$ of $B_{t}^{\mathrm{mlp2}}$ are used.
[…]
The knapsack formulation for a single stage problem of ViT is described by the following ILP-model, where all variables are binary.
  • (11)
    Maximization of the total importance score of the chosen ViT architecture.
  • (11)
    Satisfaction of constraint $C$ (i.e., MACs).
  • (12)
[…]
  • (13)
[…]
  • (14)
[…]
  • (15)
    If a neuron $f_{i}$ is picked, then its number of input weight connections is equal to the number of chosen heads convolutional filters $x_{0}$.
  • (16)
[…]
  • (17)
    The number of heads convolutional filters $N_{0}$ are chosen to be contiguous in memory.
  • (18)
[…]
  • (19)
[…]
  • (20)
[…]
$\max$ $\sum_{i=1}^{N_{o}}\Bigl{(}y_{i}\cdot P_{\mathrm{pat}_{i}}+\sum_{t=1}^{d}y_{i}\cdot\Bigl{(}P_{Q_{i}}^{t}+P_{V_{i}}+P_{K_{i}}^{t}+P_{\mathrm{proj}_{i}}^{t}\Bigr{)}\Bigr{)}$ (11)
$\quad+\sum_{t=1}^{d}\sum_{i=1}^{N_{o}}\sum_{k=1}^{N_{i}}f_{ik}^{t}\;\cdot\;P_{ik}^{\mathrm{mlp1}_{t}}$
$\quad+\sum_{t=1}^{d}\sum_{i=1}^{N_{o}}\sum_{k=1}^{N_{i}}g_{ki}^{t}\;\cdot\;P_{ki}^{\mathrm{mlp2}_{t}}$
s.t. $\sum_{i=1}^{N_{o}}\Bigl{(}y_{i}\cdot W_{\mathrm{pat}_{i}}+\sum_{t=1}^{d}y_{i}\cdot\Bigl{(}W_{Q_{i}}^{t}+W_{V_{i}}^{t}+W_{K_{i}}^{t}+W_{\mathrm{proj}_{i}}^{t}\Bigr{)}\Bigr{)}$
$\quad+\sum_{t=1}^{d}\sum_{i=1}^{N_{o}}\sum_{k=1}^{N_{i}}f_{ik}^{t}\;\cdot\;W_{ik}^{\mathrm{mlp1}_{t}}$
$\quad+\sum_{t=1}^{d}\sum_{i=1}^{N_{o}}\sum_{k=1}^{N_{i}}g_{ki}^{t}\;\cdot\;W_{ki}^{\mathrm{mlp2}_{t}}\;\leq\;C$
$\sum_{i=1}^{N_{o}}y_{i}=x_{0}\;\;\;\;\text{and}\;\;\;\;\sum_{i=1}^{N_{i}}f_{i1}^{t}=n_{\mathrm{mlp1}_{t}}\enspace\forall t$ (12)
$\sum_{k=1}^{N_{o}}f_{1k}^{t}=x_{0}\enspace\forall t$ (13)
$\sum_{k=1}^{N_{i}}g_{1k}^{t}=n_{\mathrm{mlp1}_{t}}\enspace\forall t$ (14)
$\sum_{i=1}^{N_{o}}f_{ik}^{t}\geq x_{0}-N_{o}\cdot(1-f_{1k}^{t})\enspace\forall k,t$ (15)
$\sum_{i=1}^{N_{o}}g_{ik}^{t}\geq x_{0}-N_{o}\cdot(1-g_{1k}^{t})\enspace\forall k,t$ (16)
$y_{i}\geq y_{i+1}\enspace\forall i$ (17)
$f_{ik}^{t}\leq y_{i}\;\;\;\;\;\;\;\;\;\;\text{and}\;\;\;\;g_{ik}^{t}\leq y_{i}\;\enspace\quad\quad\forall i,k,t$ (18)
$f_{ki}^{t}\geq f_{i,k+1}^{t}\;\;\;\;\text{and}\;\;\;\;g_{ik}^{t}\geq g_{i,k+1}^{t}\quad\forall i,k,t$ (19)
$f_{ik}^{t}\geq f_{i+1,k}^{t}\;\;\;\;\text{and}\;\;\;\;g_{ik}^{t}\geq g_{i+1,k}^{t}\;\quad\forall i,k,t$ (20)
3.4.4 Peak Memory Usage Constraint
Formulating structural pruning of a neural network as an iterative knapsack problem allows to add constraints to each subnetwork’s architecture by assigning different computational costs to each item in the integer program. We show how to constrain peak memory usage by limiting the summed size of the output tensors for each layer (Liberis and Lane, 2023). Given the previously defined single-stage knapsack formulation (2), $OutMem_{il}$ as the size of the output tensor of item $i$ from layer $l$, and $C_{PeakMem}$ as the maximum activation map size, the peak memory usage constraint is defined as follows:
$\sum_{i=1}^{u_{l}}x_{il}\cdot OutMem_{il}\leq C_{PeakMem}\enspace\forall l$ (21)
3.4.5 Knapsack for Language and Multimodal Models
[…]
3.4.6 Knapsack for Semantic Segmentation and Object Detection Models
REDS knapsack formulation is agnostic to the architecture of the input model, provided that the model and its loss function are differentiable, thus allowing gradient computation, required when employing gradient-based importance scores for the units. Given a pre-trained segmentation or object-detection model, REDS computes unit-wise importance scores using the respective task-specific loss (e.g., pixel-wise cross-entropy for segmentation or the combined classification and localization loss for detection) and then solves the iterative knapsack problem under specified computational or memory budget constraints given as input. The selected subnetworks can subsequently be fine-tuned, using the same task-specific loss, to ensure high predictive performance on the target task.
@@ 3.5 REDS fine-tuning
Fine-tuning models after pruning is a common practice to recover model accuracy. The REDS subnetworks found by the heuristics algorithm are fine-tuned to recover their full accuracy. For the DS-CNN, CNN and DNN architectures we first simultaneously fine-tune the encoder’s weights followed by the batch normalization layers (Qu et al., 2022). For the ViT architecture we employed a stochastic training approach following (Haberer et al., 2024) to finetune each subnetwork. For each subnetwork we slice each layer to construct a subnetwork architecture by storing only the slicing point corresponding to the number of active computational units. During fine-tuning, the weights of each subnetwork $i$ are reused by all lower-capacity subnetworks $\{i+k\}_{k=1}^{N}$ by dynamically creating a tensor slice for each layer, i.e., a tensor view object that points to the original weight tensor. For the DS-CNN, CNN and DNN architectures we balance the contribution of the loss of each individual model with parameters $\{\pi_{i}\}_{i=1}^{N}$ equal to the percentage of weights used in the encoder part of the model by the subnetwork $i$ (Qu et al., 2022).
Figure 3: Comparison of REDS with Slimmable Subnets (Yu et al., 2019) and width pruning methods (i.e., SNIP (Lee et al., 2018) and magnitude pruning) for subnetworks derived from pre-trained DS-CNN models of sizes S and L, evaluated on the Google Speech Commands dataset. The two plots on the left depict test accuracy versus inference time for each subnetwork on the Arduino Nano 33 BLE IoT platform, while the two plots on the right show test accuracy versus model size. REDS consistently outperforms Slimmable Subnets, SNIP, and magnitude pruning, achieving superior accuracy–inference time and accuracy–size trade-offs.
4 REDS Performance
4.1 Evaluation setup
We evaluate the performance of REDS, which comprises four nested subnetworks obtained by multiplying the full model MACs (100%) by the predefined constraints percentages (25%, 50%, and 75%), on three datasets: Google Speech Commands (Warden, 2018) (GSC), Fashion-MNIST (Xiao et al., 2017) and CIFAR-10 (Krizhevsky et al., 2009). In addition, we evaluate four subnetworks on GSC whose MACs budgets correspond to width-pruning ratios of 0.25, 0.5 and 0.75. Moreover, on the Visual Wake Words (VWW) dataset (Chowdhery et al., 2019), we evaluate two nested subnetworks under constraints of 50% of MACs and 250 KB peak memory usage. We also evaluate three nested subnetworks on the ImageNet-1K dataset (Deng et al., 2009). We measure peak memory usage using the open-source tflite-tools (Liberis et al., 2021), assuming that the temporary buffer space serves both input and output buffers (Chowdhery et al., 2019). MACs counts are obtained via the tfprof Tensorflow profiler.
REDS is evaluated on: three pre-trained GSC architectures (DNN, CNN, and DS-CNN) of two sizes (S and L) each introduced in Zhang et al. (2017), an ImageNet-1K (Deng et al., 2009) pretrained MobileNetV1 (Howard et al., 2017) architecture fine-tuned on the VWW dataset; and a ViT-Base (Dosovitskiy et al., 2020) model trained on ImageNet-1K. For CIFAR10 and Fashion-MNIST, we test REDS on two pre-trained, size S DS-CNN architectures. The DNN comprises a two layers fully-connected architecture with layer widths of 144 (S) and 436 (L). The CNN comprises an architecture with two convolutional layers followed by three fully-connected layers. For the size S model, the convolutional layers have widths of 28 and 30 for the S architecture; for size L, widths of 60 and 76. Both DS-CNN and MobileNetV1 comprise a standard convolutional layer followed by several depth-wise and point-wise convolutional blocks (see Fig. 2). DS-CNN has four blocks in the size S network and five in the size L, with layer widths of 64 and 276 respectively. MobileNetV1 includes 13 blocks, beginning with an initial layer width of 32 that progressively increases to 1024 with the model’s depth. ViT-Base consists of a patch-embedding layer followed by 12 transformer blocks, each comprising a multi-head self-attention sub-layer with 12 heads and a multilayer perceptron sub-layer with a hidden dimension of 3072 neurons. Each convolutional layer in the CNN, DS-CNN, and MobileNetV1 architectures is followed by batch normalization (Ioffe and Szegedy, 2015). In ViT-Base, each multi-head self-attention and multi-layer perceptron sub-layer is preceded by layer normalization (Ba et al., 2016).
Table 1 summarizes the hyper-parameters used to fine-tune the different DS-CNN networks taken from Zhang et al. (2017). For ViT we followed the hyperparameters used in Haberer et al. (2024). We used the TensorFlow (Abadi et al., 2016) version 2.12. All models were trained on a workstation with 16 NVIDIA Tesla K80 GPUs and a workstation with 60 NVIDIA A100 GPUs and 564 AMD Ryzen 3 CPUs. To conduct all our experiments and compute the baselines, we trained and optimized over 100 models.
Figure 4: Peak memory usage analysis of MobileNetV1 processed by the REDS knapsack BU heuristic. Placing constraints on MACs (left), and on both MACs and peak memory usage (right).
Table 1: Training hyper-parameters.
Hyper-parameter DNN (S/L) CNN (S/L) DS-CNN (S/L)
Batch size 100
Training epochs 75 75 250
Loss function Cross-entropy
Optimizer Adam
LR scheduler PiecewiseConstantDecay
LR $10^{-3}$, $10^{-4}$ $5\cdot 10^{-4}$, $10^{-4}$, $2\cdot 10^{-5}$ $10^{-3}$, $10^{-4}$
LR steps $1.5\cdot 10^{4}$, $2\cdot 10^{4}$, $10^{4}$, $10^{4}$ $10^{4}$, $10^{4}$, $10^{4}$
$3\cdot 10^{3}$
4.2 REDS empirical evaluation
The main body of the paper presents the results for DS-CNN architectures (S and L) on GSC, CIFAR10, Fashion-MNIST, MobileNetV1 (0.25x) on VWW and ViT on ImageNet-1K. The additional results for other architectures are available in the supplementary extended appendix. The REDS structure for the considered datasets used in the main paper comprises only four nested subnetworks; however a larger number of subnetworks does not degrade the accuracy of REDS fine-tuned on GSC (see Fig. 6). The results in Fig. 5 show REDS performance using DS-CNN architecture of size S on Fashion-MNIST and CIFAR10. BU heuristic was used to obtain the results. REDS supports a different data domain without degrading the accuracy of the pre-trained model, reported in the header row. Curiously, jointly training the subnetworks improves the performance of the pre-trained 100% MACs subnetwork, similarly as reported in Yu et al. (2019).
Table 2: Test set accuracy [%] of training S and L depth-wise separable convolutional architectures (DS-CNN) from Zhang et al. (2017): Training a network of each size from scratch ("Scratch"), conversion from a pre-trained network using "Knapsack BU " and "Knapsack TD", and training REDS structure from scratch ("REDS training"). Reported results from three independent runs. The accuracy of each 100 % network reported in Zhang et al. (2017) is listed in the header row.